# Prediction Sets and Conformal Inference with

### v3

@@ Abstract
Given data on a scalar random variable $Y$, a prediction set for $Y$ with miscoverage level $\alpha$ is a set of values for $Y$ that contains a randomly drawn $Y$ with probability $1-\alpha$, where $\alpha\in(0,1)$. Among all prediction sets that satisfy this coverage property, the oracle prediction set is the one with the smallest volume. This paper provides estimation methods of such prediction sets given observed conditioning covariates when $Y$ is censored or measured in intervals. We first characterise the oracle prediction set under interval censoring and develop consistent estimators for the oracle prediction intervals and prediction sets consisting of multiple disjoint intervals. We use conformal inference to construct a prediction set that achieves finite-sample validity under censoring and maintains consistency as sample size increases, using a conformity score function designed for interval data. The procedure accommodates the prediction uncertainty that is irreducible (due to the stochastic nature of outcomes), the modelling uncertainty due to partial identification and also sampling uncertainty that gets reduced as samples get larger. We conduct a set of Monte Carlo simulations and an application to data from the Current Population Survey. The results highlight the robustness and efficiency of the proposed methods.
Keywords: Oracle predictions sets, partial identification, nonparametric estimation, uniform consistency of set estimator.
@@ 1 Introduction
Interval data is pervasive. Surveys usually employ “bracketing” strategies to address item non-response. For example, respondents are often asked whether their money in savings accounts, in investments or their income fall in a sequence of brackets such as [\$$[\$10K,\$20K],[\$20K,\$50K]$ etc., while also given the option to provide a point answer to such questions (see Heeringa and Suzman, (1995) and Moore and Loomis, (2001)). Interval valued data also accommodate often found patterns such as right censoring, where an outcome variable is greater than a specified value (i.e., $Y\in[Y^{L},\infty)$)); left censored data, where an outcome variable is less than a certain value (i.e., $Y\in(-\infty,Y^{U}])$); and data from competing risks, where one typically observes the $\max$ or $\min$ of variables of interest. Such data typically involve outcomes of interest such as income, wealth, unemployment duration and cause-specific mortality. This paper provides inference procedures for the prediction of a variable of interest $Y$ when $Y$ is not observed, but we instead observe a sample of $[Y^{L},Y^{U}]$ such that $Y\in[Y^{L},Y^{U}]$.
More specifically, suppose we have a random variable $Y$ defined on $\mathcal{Y}\subset\mathbb{R}$ with distribution $P$. One parameter of interest is a prediction set $C$ defined as the set of $y\in\mathcal{Y}$ such that $P(y\in C)\geq 1-\alpha$ for $\alpha\in(0,1)$ where we say that such a set $C$ has miscoverage level $\alpha\in(0,1).$ The oracle prediction set is the set with the smallest volume that has this (mis)coverage property and is related to the level set of the density of $Y$ (when this density exists). With covariates, we have access to a sample $(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y}$, $i=1,\dots,n$, with supports $\mathcal{X}\subset\mathbb{R}^{d}$ and $\mathcal{Y}\subset\mathbb{R}$. We use these data to construct prediction sets $C_{n}(x)\equiv C(x;X_{1},Y_{1},\dots,X_{n},Y_{n})\subset\mathcal{Y}$ for a given miscoverage level $\alpha$ at $x\in\mathcal{X}$. It is desirable that such prediction set $C_{n}$ is both consistent for the oracle prediction set and has a finite sample coverage property:
@@
where $\operatorname{P}$ is the distribution over $(X_{i},Y_{i})$ for $i=1,\dots,n+1$. This paper constructs such a set when $Y$ is interval valued. The exact finite-sample coverage property that our construction provides is given in more details in Section 4 .
[…]
@@
This “operational density” summarizes both the sampling uncertainty in estimating $\hat{\beta}$ and the disturbance uncertainty that comes from the intrinsic randomness in the conditional distribution $Y_{n+1}\mid X_{n+1}$. One can use it to derive a forecast density with a particular coverage property.
@@
The prediction set is closely related to quantile methods: the minimal volume prediction set would be given by $C(X_{n+1})\equiv[\tau_{0}(X_{n+1}),\tau_{1}(X_{n+1})]$, where $\tau_{0}(X_{n+1})$ and $\tau_{1}(X_{n+1})$ are appropriately chosen quantiles for $Y$ given $X$ so that one can obtain a coverage rate $(1-\alpha)$ at a minimum volume. Taking the Gaussian specification above, for example, with $\alpha=5\%$ and assuming $\beta$ to be known, $\tau_{0}(X_{n+1})=X_{n+1}^{\prime}\beta-1.96\sigma$, $\tau_{1}(X_{n+1})=X_{n+1}^{\prime}\beta+1.96\sigma$ where $1.96\sigma$ is the $95\%$-quantile for $|\varepsilon_{n+1}|\equiv|Y_{n+1}-X_{n+1}^{\prime}\beta|$. When $\sigma$ is known as presumed above, but $\beta$ is not, estimates for these can be obtained as $\hat{\tau}_{0}(X_{n+1})=X_{n+1}^{\prime}\hat{\beta}-1.96\sigma$ and $\hat{\tau}_{1}(X_{n+1})=X_{n+1}^{\prime}\hat{\beta}+1.96\sigma$, where $\hat{\beta}$ is the OLS or ML estimate for $\beta$. This nonetheless ignores the parameter uncertainty in estimating $\beta$. This is not a problem if we are just interested in interval consistency, but as noted in Romano et al., (2019), replacing $\beta$ with $\hat{\beta}$ “is not guaranteed to satisfy the coverage statement (…) when $C(X_{n+1})$ is replaced by the estimated interval $\hat{C}(X_{n+1})$.” To address this shortcoming, Romano et al., (2019) employ conformal inference and modify the interval $\hat{C}(X_{n+1})=[\hat{\tau}_{0}(X_{n+1}),\hat{\tau}_{1}(X_{n+1})]$ by adjusting the endpoints to $\hat{\tau}_{0}(X_{n+1})-\vartheta_{1-\alpha}$ and $\hat{\tau}_{1}(X_{n+1})+\vartheta_{1-\alpha}$. Here, $\vartheta_{1-\alpha}$ is the $(1-\alpha)(1+1/n_{2})$-th empirical quantile of the set of conformity scores $\left\{\max\{\hat{\tau}_{0}(X_{i})-Y_{i},Y_{i}+\hat{\tau}_{1}(X_{i})\}:i\in\mathcal{I}_{2}\right\}$ estimated in a calibration sub-sample $\mathcal{I}_{2}$ with sample size $n_{2}$. This corrects for under- or over-coverage and allows for the construction of prediction intervals satisfying desired coverage levels in finite samples using quantile estimates. Our method generalises this idea to interval-valued outcomes. Further details on the conformal inference procedure are provided in Section 4. We also note that although we can construct valid prediction interval with quantile estimates, such intervals are not necessarily minimal in volume, and it is not applicable the oracle prediction set consists of multiple disjoint intervals. Our method addresses these limitations and offers a complementary perspective to traditional quantile-based methods.
When $Y$ is interval-valued and without parametric assumptions, we use recent developments on conformal inference to show formally that a particularly constructed predictive set $\tilde{C}_{n}$ has finite sample coverage property while also enjoying the asymptotic property of being consistent for the oracle prediction set under the partially identified true population distribution. Our conformal procedure designed when $Y$ is interval-valued in turn requires the definition of particular score functions to handle this interval censoring. This score is similar to ones used in the conformal literature but is here modified to handle censoring. In particular, the score function we use is negative when the constructed prediction set contains the interval outcomes, but is otherwise positive, and as a result accounts for both over- and under-coverage as in Romano et al., (2019).
@@
In order to obtain the proposed conformal prediction set, we first characterise the oracle prediction sets under interval censoring and provide a feasible optimisation problem that relates the oracle prediction sets to the conditional distribution of the observed upper and lower brackets, $(Y^{L},Y^{U})$. We use a kernel based conditional distribution function estimator as a basis for the construction of a consistent set estimator for the oracle prediction set. The oracle prediction set may be the union of a collection of disjoint intervals in cases for instance where the prediction density is multimodal. We provide an approach that allows for the set estimator to be a union of disjoint intervals with the prescribed coverage. In cases with a multimodal density, a union of intervals can have a smaller volume than one interval.
@@
We then adapt recent work on conformal inference to propose a conformal prediction set using a sample of exchangeable data. This proposed estimator is shown to be both consistent and have guarantees for finite sample coverage. As a competing alternative prediction set estimator, we show in the appendix that a conformal quantile regression method provides a valid prediction set for interval valued outcomes which can be easier to implement in practice, but is in general not efficient in terms of the volume of the prediction set. In addition, quantile regression based methods are not suitable when the prediction set is not a single interval.
Interval censoring may arise from fixed bracketing and random censoring from “unfolding brackets” or survey schemes, see Heeringa and Suzman, (1995) and Moore and Loomis, (2001). See also the bracketing approach to outcomes using the Health and Retirement Survey in Manski and Tamer, (2002). We thus illustrate our construction using a set of Monte Carlo experiments, where we consider both fixed and random censoring. Our proposed prediction set is shown to have adequate finite sample coverage properties and to have smaller volume compared with the alternative prediction sets constructed with quantile regression methods. In addition, we construct prediction sets using the Current Population Survey (CPS) data. Among people who report their income, a significant fraction (more than 20%) of the observations are interval valued. Interval valued observations are sometimes discarded or handled by imputation methods that rely on assumptions such as missing at random. Our proposed prediction set is robust in the sense that we do not require a specification of the censoring mechanism and we demonstrate using a hold-out subset of the CPS data that our prediction set has better finite sample coverage of the interval valued incomes compared with prediction sets constructed with hot-deck imputation in place. Naturally, without relying on imputation algorithms, our oracle prediction sets are wider than ones obtained with uncensored outcomes, but our constructed sets are more robust and hence more trustworthy.
@@
The paper combines insights from the partial identification literature and conformal inference, which has recently been gaining significant interest. The partial identification literature (see Manski, (2003)) takes the view that one should try and make inferences under minimal assumptions, even if these assumptions do not allow for point identification of the parameter of interest. The motivation for such a view is that inferences are less credible and conclusions are suspect when these rely on non-testable or non credible assumptions. In our context for example, if we are interested in $\beta=E[Y]$ and we observe $[Y^{L},Y^{U}]$ rather than $Y$, then a partial identification approach would just say that, given the data, the identified set for $\beta$ (or all we can learn about $\beta$ given the data) is the interval $\left[E[Y_{0}],E[Y_{1}]\right].$ Imputation based inference uses models to predict a $Y^{P}$ within $[Y^{L},Y^{U}]$ and takes $E[Y^{P}]$ as a proxy for $\beta=E[Y]$. The credibility of this exercise depends on the imputation model, i.e. the set of assumptions it relies on. In the context of interval data, Manski and Tamer, (2002) study identification of linear regression coefficients when the data (on regressors or outcomes) is interval valued using a partial identification approach. See Tamer, (2010), Molinari, (2020), and Kline and Tamer, (2023) for reviews on the developments in the partial identification literature.
@@
In order to obtain a prediction interval that has finite-sample coverage guarantees, we use conformal inference, which is a distribution-free method that builds upon the idea of exchangeability and permutation tests (Vovk et al., (2005), Vovk et al., (2009)). The idea of conformal inference has become increasingly popular in the machine learning literature since it requires no assumptions on the underlying distribution other than the observations being exchangeable and it also allows the use of any prediction protocol that treats the data symmetrically and hence preserves exchangeability. (Barber et al., (2023) discusses the validity of conformal inference when the data is not i.i.d. and the predictor is not symmetric.) Conformal inference has been applied to obtain valid distribution-free prediction sets (Lei et al.,, 2013), in non-parametric regression (Lei and Wasserman,, 2014) and in high-dimensional settings (Lei et al.,, 2018). Romano et al., (2019) considers conformal quantile regression to construct prediction intervals instead of conditional mean regression. Sadinle et al., (2019) proposes a way to construct set-valued classifiers. Chernozhukov et al., (2021) considers construction of valid prediction intervals with distributional regression method. We add to this literature by allowing the outcome observations to be interval-censored or interval-valued. Relatedly,Candès et al., (2023) considers conformal inference for survival analysis. Using our notation, this amounts to observing either $Y_{i}$ or the interval $(Y_{i}^{L},\infty)$ for each individual $i$. They propose a weighted conformal inference procedure (Tibshirani et al.,, 2019) under the assumption that censoring is conditionally independent of her outcome given predictors $X_{i}$. We take a different approach and ask how much information we can extract from a sample with interval-valued outcomes. For example, our procedure allows for the case when all outcomes are interval censored (i.e., $Y_{i}$ is never observed) and the procedure uses the information from the upper bracket $Y^{U}_{i}$ when it is not $\infty$. Also, our procedure does not require the conditional independence assumption which can be violated in many empirical contexts.
@@
There are several related areas connected to our theoretical development. The classical way to construct estimators for the oracle prediction set relies on the level set estimation, (see, for example, Wilks, (1941), Polonik, (1995) and Samworth and Wand, (2010)). The uniform consistency of the nonparametric estimator of the conditional distribution is an extension of classical results on the uniform convergence rate of the kernel-based density and regression estimators such as Einmahl and Mason, (2000), Gine, (2002) and Hansen, (2008). This paper is also related to the literature on quantile regression (see Koenker, (2017) for a review).
@@
This paper is organised as follows. In Section 2, we provide the setup and definitions of optimal sets and approaches to obtaining estimators of such sets with censoring. Section 3 provides a study of consistency of such sets using kernel estimators of the conditional distribution. Section 4 provides results on finite sample coverage using a conformal inference procedure under interval censoring. Section 5 provides some Monte Carlo evidence and Section 6 applies our methods to US CPS Data. Section 7 concludes. The proofs for the theorems in the main text as well as some further results and comments are collected in the Appendix.

### v4

# Weiguang Liu UCL Áureo de Paula

@@ Abstract
Given data on a random variable $Y$, a prediction set with miscoverage level $\alpha\in(0,1)$ is a set that contains a new draw of $Y$ with probability $1-\alpha$. Among all prediction sets satisfying this coverage property, the oracle prediction set is the one with minimal volume. The oracle prediction set offers a complementary view of the distribution of $Y$, beyond point estimators such as the mean and quantiles, and has attracted considerable interest recently. This paper develops methods for estimating such prediction sets conditional on observed covariates when $Y$ is censored or interval-valued. We characterise the oracle prediction set under partial identification induced by interval censoring and propose consistent estimators for both oracle prediction intervals and more general oracle prediction sets consisting of multiple disjoint intervals. In addition, we apply conformal inference to construct finite-sample valid prediction sets for interval outcomes that remain consistent as the sample size grows, using a conformity score tailored to interval data. The proposed procedure accounts for irreducible prediction uncertainty due to the stochastic nature of outcomes, modelling uncertainty arising from partial identification, and sampling uncertainty that vanishes as sample size increases. We conduct Monte Carlo simulations and two empirical applications using UK job postings data and the US Current Population Survey. The results demonstrate the robustness and efficiency of the proposed methods.
Keywords: Prediction sets; Partial identification; Nonparametric estimation; Conformal Prediction.
@@ 1 Introduction
Interval data are pervasive. Surveys often resort to “bracketing” strategies to address item non-response. For example, respondents are often asked whether their money in savings accounts, in investments, or their income falls into a sequence of brackets such as $\mathdollar 10K,\mathdollar 20K]$, $\mathdollar 20K,\mathdollar 50K]$, etc, while also being given the option to provide a point answer to such questions (see Heeringa and Suzman (1995) and Moore and Loomis (2001)). Interval-valued data also accommodate common patterns such as right censoring, where an outcome variable is greater than a specified value (i.e., $Y\in[Y^{L},\infty)$); left-censoring, where an outcome variable is less than a certain value (i.e., $Y\in(-\infty,Y^{U}]$); and data from competing risks, where one typically observes the $\max$ or $\min$ of variables of interest. Such data typically involves outcomes of interest such as income, wealth, unemployment duration, and so on. This paper provides a new method for predicting an outcome variable of interest $Y$ when we observe a sample of intervals $Y^{L},Y^{U}]$ such that $Y\in[Y^{L},Y^{U}]$ rather than direct observations of $Y$.
More specifically, suppose we have a random variable $Y$ defined on $\mathcal{Y}\subset\mathbb{R}$ with distribution $P_{Y}$. One parameter of interest is a prediction set $C\subset\mathcal{Y}$ with miscoverage level $\alpha\in(0,1)$, which is any measurable subset of $\mathcal{Y}$ such that $P_{Y}(Y\in C)\geq 1-\alpha$. The oracle prediction set is the set with the minimal volume that has this coverage property, which is related to the level set of the density of $Y$ when this density exists. With covariates, we have access to a sample $(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y}$, $i=1,\dots,n$, with supports $\mathcal{X}\subset\mathbb{R}^{d}$ and $\mathcal{Y}\subset\mathbb{R}$. We use these data to construct prediction sets $C_{n}(x)\equiv C(x;X_{1},Y_{1},\dots,X_{n},Y_{n})\subset\mathcal{Y}$ for a given miscoverage level $\alpha$ at $x\in\mathcal{X}$. It is desirable that such a prediction set $C_{n}$ is both consistent for the oracle prediction set and has a finite-sample coverage property:
@@
where $\operatorname{P}$ is the distribution over $(X_{i},Y_{i})$ for $i=1,\dots,n+1$. This paper constructs such a set when $Y$ is interval-valued. The exact finite-sample coverage property that our construction provides is given in more detail in Section 4.
Such an “oracle prediction” exercise is familiar in the literature, especially in the context of forecasting in linear regressions (see Diebold (2015)). For example, assuming that the outcome $Y$ satisfies $Y_{i}\mid X_{i}\sim\mathcal{N}(X_{i}^{\prime}\beta,\sigma^{2})$ with known $\sigma^{2}$, the population oracle prediction set is $X_{n+1}^{\prime}\beta\pm 1.96\sigma$. Given data $\mathcal{D}=\left\{(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})\right\}$, a consistent estimator for such a set is $X_{n+1}^{\prime}\hat{\beta}\,\pm\,1.96\sigma$, where $\hat{\beta}$ is the least squares estimator of $\beta$. On the other hand, “an operational density forecast that accounts for parameter uncertainty” is (Diebold (2015)):
@@
This “operational density” summarises both the sampling uncertainty in estimating $\hat{\beta}$ and the disturbance uncertainty arising from the intrinsic randomness in the conditional distribution $Y_{n+1}\mid X_{n+1}$. One can use it to derive a forecast density with a particular coverage property.
@@
The prediction set is closely related to quantile methods if we have direct observations of the outcome variable $Y$. If the oracle prediction set is a single interval, it is given by $C(X_{n+1})\equiv[\tau_{0}(X_{n+1}),\tau_{1}(X_{n+1})]$, where $\tau_{0}(X_{n+1})$ and $\tau_{1}(X_{n+1})$ are appropriately chosen quantiles for $Y$ given $X$ so that one can obtain a coverage rate $(1-\alpha)$ with minimal volume. Taking the Gaussian specification above, with $\alpha=5\%$ and assuming $\beta$ is known, $\tau_{0}(X_{n+1})=X_{n+1}^{\prime}\beta-1.96\sigma$ and $\tau_{1}(X_{n+1})=X_{n+1}^{\prime}\beta+1.96\sigma$, where $1.96\sigma$ is the $95\%$-quantile of $|\varepsilon_{n+1}|\equiv|Y_{n+1}-X_{n+1}^{\prime}\beta|$. When $\sigma$ is known as presumed above, but $\beta$ is not, estimates for these can be obtained as $\hat{\tau}_{0}(X_{n+1})=X_{n+1}^{\prime}\hat{\beta}-1.96\sigma$ and $\hat{\tau}_{1}(X_{n+1})=X_{n+1}^{\prime}\hat{\beta}+1.96\sigma$, where $\hat{\beta}$ is the OLS or ML estimate of $\beta$. This nonetheless ignores the parameter uncertainty in estimating $\beta$. This is not a problem if we are just interested in interval consistency, but as noted in Romano et al. (2019), replacing $\beta$ with $\hat{\beta}$ “is not guaranteed to satisfy the coverage statement (…) when $C(X_{n+1})$ is replaced by the estimated interval $\hat{C}(X_{n+1})$.” To address this shortcoming, Romano et al. (2019) employ conformal inference and modify the interval $\hat{C}(X_{n+1})=[\hat{\tau}_{0}(X_{n+1}),\hat{\tau}_{1}(X_{n+1})]$ by adjusting the endpoints to $\hat{\tau}_{0}(X_{n+1})-\vartheta_{1-\alpha}$ and $\hat{\tau}_{1}(X_{n+1})+\vartheta_{1-\alpha}$. Here, $\vartheta_{1-\alpha}$ is the $(1-\alpha)(1+1/n_{2})$-th empirical quantile of the set of conformity scores $\left\{\max\{\hat{\tau}_{0}(X_{i})-Y_{i},Y_{i}-\hat{\tau}_{1}(X_{i})\}:i\in\mathcal{I}_{2}\right\}$ estimated in a calibration sub-sample $\mathcal{I}_{2}$ with sample size $n_{2}$. This corrects for under- or over-coverage and allows for the construction of prediction intervals satisfying desired coverage levels in finite samples using quantile estimates.
However, several questions and limitations surface when the outcome variable $Y$ is interval-censored. First, the definition of the oracle prediction set needs to be extended in partially identified models. We also note that although we can construct a valid prediction interval using quantile estimates, such intervals are not necessarily minimal in volume without assumptions on the shape of the distribution (such as the Gaussian assumption in the example above), and this approach is not applicable when the oracle prediction set consists of multiple disjoint intervals. Furthermore, a new conformity score is needed to accommodate interval-valued observations. Our method addresses these limitations and offers a complementary perspective on the distribution of outcomes beyond existing quantile-based methods.
@@
We first define the oracle prediction sets under partial identification. We then characterise the oracle prediction sets under interval censoring via a feasible optimisation problem that relates the oracle prediction sets to the conditional distribution of the observed upper and lower brackets, $(Y^{L},Y^{U})$. We use a kernel-based conditional distribution function estimator as a basis for the construction of a consistent set estimator for the oracle prediction set. The oracle prediction set may be the union of disjoint intervals, for instance, when the prediction density is multimodal. We provide an approach that allows for the prediction set estimator to be a union of disjoint intervals with the prescribed coverage. In cases with a multimodal density, a union of intervals can have a smaller volume than one interval. We also allow for both random censoring with continuously distributed brackets and fixed censoring with discretely distributed brackets, such as those arising from “unfolding brackets” or specific survey schemes, see Heeringa and Suzman (1995) and Moore and Loomis (2001). See also the bracketing approach to outcomes using the Health and Retirement Survey in Manski and Tamer (2002).
@@
We use recent developments on conformal inference to propose a conformal prediction set $\tilde{C}_{n}$ using a sample of exchangeable data, which has finite-sample validity while also enjoying the asymptotic property of being consistent for the oracle prediction set under partial identification. Our conformal procedure is designed for interval-valued outcomes, which requires the design of a particular conformity score function to handle interval data. In particular, the score function is non-positive if and only if the constructed prediction set contains the interval outcome; this conformity score function therefore accounts for both over- and under-coverage, as in Romano et al. (2019). As a competing alternative prediction set estimator, we show in the appendix that a conformal quantile regression method provides a valid prediction interval for interval-valued outcomes, which can be easier to implement in practice, but in general will not be efficient in terms of the volume of the prediction set. In addition, quantile regression-based methods are not suitable when the prediction set is not a single interval.
We illustrate our construction using a set of Monte Carlo experiments, where we consider both fixed and random censoring. Our proposed prediction set is shown to have robust finite sample coverage properties and to have a smaller volume compared with the alternative prediction sets constructed with quantile regression methods.
We apply our methods to two real-world datasets. The first is the UK job advert dataset from Adzuna, an online UK job advert aggregator. Adzuna data are widely used as an indicator of UK economic activity. One feature of these data is that many job ads report wages as ranges. We use these data to construct predictive wage intervals as a function of UK location and job categories. The second application uses the Current Population Survey (CPS) data. Among people who report their income, a significant fraction of the observations are interval-valued. Interval-valued observations are sometimes discarded or handled by imputation methods that rely on assumptions such as missing at random and may cause bias if misspecification happens. Our proposed prediction set is robust in that we do not require a specification of the censoring mechanism.
@@
The paper combines insights from the partial identification literature and conformal inference, which has recently been gaining significant interest. The partial identification literature (see Manski (2003)) takes the view that one should try to make inferences under minimal assumptions, even if these assumptions do not allow for point identification of the parameter of interest. The motivation for such a view is that inferences are less credible and conclusions are suspect when they rely on non-testable or non-credible assumptions. In our context for example, if we are interested in $\beta=E[Y]$ and we observe $Y^{L},Y^{U}]$ rather than $Y$, then a partial identification approach would just say that, given the data, the identified set for $\beta$ (or all we can learn about $\beta$ given the data) is the interval $\left[E[Y^{L}],E[Y^{U}]\right].$ Imputation based inference uses models to predict a $Y^{P}$ within $Y^{L},Y^{U}]$ and takes $E[Y^{P}]$ as a proxy for $\beta=E[Y]$. The credibility of this exercise depends on the imputation model, i.e. the set of assumptions it relies on. In the context of interval data, Manski and Tamer (2002) study identification of linear regression coefficients when the data (on regressors or outcomes) is interval valued using a partial identification approach. See Tamer (2010), Molinari (2020), and Kline and Tamer (2023) for reviews on the developments in the partial identification literature.
@@
In order to obtain a prediction interval that has finite-sample coverage guarantees, we use conformal inference, which is a distribution-free method that builds upon the idea of exchangeability and permutation tests (Vovk et al. (2005), Vovk et al. (2009)). The idea of conformal inference has become increasingly popular in the machine learning literature since it requires no assumptions on the underlying distribution other than the observations being exchangeable, and it also allows the use of any prediction protocol that treats the data symmetrically and hence preserves exchangeability. Barber et al. (2023) discusses the validity of conformal inference when the data is not i.i.d. and the predictor is not symmetric. Conformal inference has been applied to obtain valid distribution-free prediction sets (Lei et al., 2013), in non-parametric regression (Lei and Wasserman, 2014) and in high-dimensional settings (Lei et al., 2018). Romano et al. (2019) considers conformal quantile regression to construct prediction intervals instead of conditional mean regression. Sadinle et al. (2019) proposes a way to construct set-valued classifiers. Chernozhukov et al. (2021) considers the construction of valid prediction intervals with the distributional regression method. We add to this literature by allowing the outcome observations to be interval-censored or interval-valued. Relatedly,Candès et al. (2023) considers conformal inference for survival analysis. Using our notation, this amounts to observing either $Y_{i}$ or the interval $(Y_{i}^{L},\infty)$ for each $i$. They propose a weighted conformal inference procedure (Tibshirani et al., 2019) under the assumption that censoring is conditionally independent of her outcome given predictors $X_{i}$. We take a different approach and ask how much information we can extract from a sample with interval-valued outcomes. For example, our procedure allows for the case when all outcomes are interval censored (i.e., $Y_{i}$ is never observed) and the procedure uses the information from the upper bracket $Y^{U}_{i}$ when it is not $\infty$. Also, our procedure does not require the conditional independence assumption, which can be violated in many empirical contexts.
@@
There are several related areas connected to our theoretical development. The classical way to construct estimators for the oracle prediction set relies on the level set estimation (see, for example, Wilks (1941), Polonik (1995) and Samworth and Wand (2010)). The uniform consistency of the nonparametric estimator of the conditional distribution is an extension of classical results on the uniform convergence rate of the kernel-based density and regression estimators, such as Einmahl and Mason (2000), Gine (2002) and Hansen (2008). This paper is also related to the literature on quantile regression (see Koenker (2017) for a review).
@@
This paper is organised as follows. In Section 2, we provide the setup and definitions of optimal sets and approaches to obtaining estimators of such sets with censoring. Section 3 provides a study of the consistency of such sets using kernel estimators of the conditional distribution. Section 4 provides results on finite sample coverage using a conformal inference procedure under interval censoring. Section 5 provides some Monte Carlo evidence, and Section 6 applies our methods to UK job postings data as well as the US CPS (Current Population Survey) Data. Section 7 concludes. The proofs for the theorems in the main text, as well as some further results and comments, are collected in the Appendix.